OpenStudy (anonymous):

How do I solve |x + 9| = 18?

OpenStudy (anonymous):

subtract 9

OpenStudy (anonymous):

|x| = 9?

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

or -27 right?

OpenStudy (mertsj):

The absolute value = 18 means that the numbers in question are exactly 18 units from 0 That means that the number is either 18 or -18 and so you write: x+9=18 OR x+9=-18.

OpenStudy (anonymous):

how do you get -27?

OpenStudy (mertsj):

Solve both of those equations and you'll have your two answers.

OpenStudy (anonymous):

because 9+18 =27

OpenStudy (mertsj):

Read what I posted and you will know.

OpenStudy (anonymous):

Thanks everyone, I get it now.